# Response Tuning: Aligning Large Language Models without Instruction

### v1

@@ 5.3 Results
Figure 6: Safety evaluation results for RT and IT models trained with safety-focused examples. The results show that response space supervision alone can enable RT models to refuse unsafe queries and achieve similar refusal rates to IT models trained using safety-focused paired data. We find no significant differences in false refusal rates between the two models.
6 Embedding Behavioral Guidance in Response Space for Safety Alignment
Motivation. Similar to instruction-following alignment, recent work demonstrates that incorporating only a small set of safety-focused examples—unsafe queries paired with refusal responses—into IT data can enable models to reject unsafe inputs (Bianchi et al., 2024; Zhou et al., 2024). On the other hand, our previous experiments have shown the abilities like instruction-following are also acquired during pre-training. Based on these observations, we hypothesize that the risk assessment capabilities required to evaluate the risk of queries are also acquired during pre-training. Therefore, we investigate whether properly establishing the response space—by including refusal responses that demonstrate how to handle unsafe queries—can induce responsible behaviors even without using paired data.
@@ 6.1 Experimental Setup
Response tuning with refusals. We incorporate contextual refusals—which express an inability to assist along with a reason explaining how the model should handle unsafe queries—into the response subsets of Alpaca, Dolly, and LIMA datasets. These refusals are obtained from existing safety-focused pairs using pattern matching (Bianchi et al., 2024). To control their proportion, we add 100, 200, and 500 examples into randomly sampled sets of 1,000333We unify the size to match that of LIMA. responses from the base RT dataset.
Evaluation. We compare safety-enhanced RT models with their IT counterparts trained using the same responses but with paired instructions. We measure the Refusal Rate (RR) against unsafe queries using benchmarks such as AdvBench (Zou et al., 2023), MaliciousInstruct (Huang et al., 2024), HarmBench (Mazeika et al., 2024), and XSTest (Röttger et al., 2024). To consider exaggerated safety (over refusal) behaviors, we also measure the False Refusal Rate (FRR) using benign queries from XSTest that are lexically similar to unsafe ones (Röttger et al., 2024). Existing methods for evaluating refusals often involve significant costs, require human effort, or cannot measure FRR. Therefore, we utilize a cost-efficient, open-source LLM—Llama-3.1-70B-Instruct (Dubey et al., 2024)—to classify whether a response is a refusal or compliance, i.e., whether the response declines to assist or attempts to provide the requested information. However, our initial manual inspection found that simply instructing the model to perform this classification yields poor performance due to tricky cases:
@@
To address these issues, we designed a universal and reliable judge prompt (see Appendix A). This judge is validated using a set of 120 examples consisting of 60 compliance and 60 refusal responses, each containing 30 of the edge cases described above. Our engineered prompt, which induces chain-of-thought reasoning (Wei et al., 2022b), achieved 98.33% accuracy on this test set. We use this judge for all safety evaluations.
@@ 6.2 Results
Response supervision alone can elicit responsible behaviors. Figure 6 shows the evaluation results for the Gemma-2-9B model trained based on LIMA. The results for the other models and datasets are available in Appendix C. Notably, RT models trained with refusals exhibit substantially higher RR compared to those trained without refusals, indicating that they are able to handle unsafe queries as suggested in training responses. We also find that their FRR falls within a reasonable range. Although they require more data to achieve a safety level close to IT counterparts, their refusal rates are largely on par with IT counterparts that are additionally supervised from mappings between unsafe queries and refusals. These results indicate that complex capabilities, such as safety assessment, are largely acquired during pre-training. LLMs can generalize the guidance for handling queries embedded in the training response distribution even without supervision from explicit instruction-response mappings.
The gap in refusal rates between RT and IT models shrinks with model scale. Our experiments using Gemma-2-9B and Gemma-2-2B models reveal that the difference in safety performance between RT and IT models diminishes as the size of the base model increases. While smaller models like Gemma-2-2B show a noticeable gap—with IT models outperforming RT models in RR—larger models such as Llama-3.1-8B and Gemma-2-9B exhibit minimal differences. In some cases, the larger RT models achieve refusal rates comparable to their IT counterparts. This suggests that larger models have a greater capacity to internalize and generalize safety behaviors from limited response supervision. Please refer to Appendix C for the detailed numerical report.
@@ 7 In-Context Response Learning
Motivation. We further validate our hypothesis—that establishing an appropriate output space alone can enable LLMs to behave as desired chat assistants—in an in-context learning setting. To this end, we test whether untuned base LLMs can helpfully and safely respond to user queries given only response demonstrations.
Experimental setup. We remove instruction-response mappings from URIAL (Lin et al., 2024), which consists of 4 instruction-response pairs including one pair of unsafe instruction and refusal. We refer to this new version as URIAL-R. We then evaluate these models using two different base LLMs, Llama-3.1-8B and Gemma-2-9B, with the JustEval benchmark. We employ greedy decoding with a maximum generation length of 2,048 tokens. To further elucidate the importance of response supervision, we include zero-shot templated prompting (Lin et al., 2024) as a baseline in our evaluation. The prompts and details of the setup can be found in Appendix A.
(f) Base LLM: Llama-3.1-8B (Dubey et al., 2024)
(g) Base LLM: Gemma-2-9B (Riviere et al., 2024)
Figure 7: GPT-4 evaluation of response quality for URIAL and URIAL-R. The test is conducted using the 1,000 test instructions from the JustEval benchmark, including the safety test set. The results show that URIAL-R achieves similar performance to URIAL across all metrics in both base models, despite not being prompted with instruction-response mappings.
Results. Figure 7 presents the evaluation results. The findings indicate that the base LLM, when provided with the URIAL-R prompt containing only response demonstrations, performs comparably to the model prompted with the original URIAL, which includes instruction-response pairs. The scores across all metrics are similar between the two prompts. Additionally, URIAL-R significantly outperforms the zero-shot prompting baselines across all metrics. These results demonstrate that response demonstrations alone can elicit helpful and safe assistant behaviors in LLMs, highlighting the impact of supervision from response distribution and latent capabilities of pre-trained LLMs. This reinforces our earlier conclusion that the capabilities required to behave as chat assistants are largely inherent in pre-trained LLMs and can be activated without explicit supervision from instruction-response mappings.
@@ 8 Conclusion
In this paper, we explore the role of establishing an output space in transforming pre-trained LLMs into instruction-following and safe assistants. We propose Response Tuning (RT), a method that focuses solely on response space supervision by eliminating the instruction-conditioning step in instruction tuning. Our extensive experiments demonstrate that RT models, trained only on responses without paired instructions, effectively respond to a wide range of user queries and responsibly handle unsafe requests by generalizing embedded behavioral guidance in the response space. These findings suggest that many of the capabilities required for alignment are already inherent in pre-trained models and can be activated by establishing a proper response space. Our work emphasizes the impact of output space supervision in alignment and highlights the potential of inherent capabilities in LLMs.
Limitations & Future Work
Our study focuses on the core objectives of LLM alignment—instruction-following and safety. Future work could explore controlling response distribution to achieve more complex alignment objectives, such as mitigating sycophancy or social bias (Perez et al., 2023; Sharma et al., 2024). Additionally, our work relies on empirical results from the ablation study; direct investigations into the inherent capabilities of pre-trained LLMs, such as extracting their semantic features (Bricken et al., 2023; Templeton et al., 2024), could further clarify the role of the alignment stage and potentially improve efficiency. Moreover, future research might adopt analytical approaches for automatic selection or fine-grained control over the response distribution to achieve better alignment.
@@
Our study involves human evaluations to evaluate instruction-following LLMs. The evaluators were hired in compliance with local laws and were paid appropriate compensation. The authors manually reviewed the LLM responses flagged by OpenAI moderation API and confirmed that these pose no harm to human evaluators. In addition, evaluators had the right to immediately stop the evaluation if they wished, and were encouraged to discuss any discomfort with the authors. This work is dedicated to understanding the LLMs safety alignment stage. While we publicly release the codes for safety evaluations, we decide not to release the refusal judge validation set to prevent potential misuse of unsafe or illegal information.
@@ Acknowledgements
We thank the members of the UNIST Language and Intelligence Lab for their helpful comments. Seokhyun An is supported by the Korea Presidential Science Scholarship. This work was supported by grants from the Institute of Information & Communications Technology Planning & Evaluation (IITP), funded by the Ministry of Science and ICT (MSIT) of the Korea government: No.RS-2020-II201336 (Artificial Intelligence Graduate School Support Program at UNIST) and IITP-2024-RS-2024-00360227 (Leading Generative AI Human Resources Development). We also acknowledge the 2022 Research Fund (1.220140.01) of UNIST.
@@ A.2 Automatic Evaluations
  • MMLU (Hendrycks et al., 2021): We use the script from the open-instruct repository (Ivison et al., 2023) for evaluation. [^6] https://github.com/allenai/open-instruct Exact-match accuracy is reported in a zero-shot setting.
@@
Safety evaluation. We evaluate the safety of models by measuring the refusal rates for unsafe instructions and false refusal rates for benign instructions using multiple safety benchmarks. For HarmBench (Mazeika et al., 2024), we report the average refusal rates for standard and contextual attack subsets. We use Llama-3.1-70B-Instruct (Dubey et al., 2024) with our judge prompt to detect refusals (see Table 9). This judge is validated using a set of 120 examples consisting of 60 compliance and 60 refusal responses, each containing 30 of the edge cases described in our experiment section. The edge cases are generated using GPT-4 and our internal jailbroken LLMs.
@@ Appendix B Experimental Setup
Response refinement. We use Llama-3.1-70B-Instruct (Dubey et al., 2024) with our refinement prompt. This prompt can be found in Table 5.
Response in-context learning. The simplified template of URIAL (Lin et al., 2024), URIAL-R, and zero-shot template prompt used for the evaluation can be found in Tables  67 and 8, respectively. We use urial1kv4 prompt in their official repository as a base URIAL prompt. [^7] https://github.com/Re-Align/URIAL/blob/main/urial_prompts/inst_1k_v4.txt.md Full version of URIAL-R prompt can be found in our code repository. The generation of LLM is truncated by the response marker of URIAL (‘‘‘).
@@ Appendix C Full Experimental Results
  • Preference evaluation results for models trained using refined responses: See Table 13.
  • Safety evaluation results: See Tables 14 and 15.
Appendix D Data Examples
Examples of training data are presented in the following tables:
  • Refined responses: See Table 16.
  • Contextual refusals: See Table 17.
Appendix E Model Output Examples
@@
  • Llama-3.1-8B (Dubey et al., 2024): See Table 18 and 19 (for the model trained using refined responses).
  • Gemma-2-9B (Riviere et al., 2024): See Table 20 and 21 (for the model trained with refusals).

### v2

@@ 4.1 Experimental Setup
Figure 6: Refusal evaluation results for RT and IT models trained on datasets that include refusal cases. The results indicate that RT models can refuse unsafe queries and achieve refusal rates approaching those of IT models trained on paired data. We find no substantial differences in false refusal rates between the two models. See Tables 11 and 12 for the results of the other models.
  • Alpaca (Taori et al., 2023): 52,000 instruction-response pairs generated using the Self-Instruct (Wang et al., 2023b) framework. We use its cleaned version, which generates responses with GPT-4 (OpenAI, 2023). [^3] https://huggingface.co/datasets/yahma/alpaca-cleaned
  • Dolly (Conover et al., 2023): 15,000 instruction-response pairs manually crafted by human annotators.
  • LIMA (Zhou et al., 2024): 1,000 instruction-response pairs curated from various sources—including web data from Stack Exchange, wikiHow, and Reddit—as well as examples manually written by the authors and sourced from Super-NaturalInstructions (Wang et al., 2022).
Training setup. We use a parameter-efficient fine-tuning method, QLoRA (Dettmers et al., 2024), which has been shown to match the performance of full 16-bit fine-tuning while significantly reducing memory footprint. LoRA adapters (Hu et al., 2022) are applied to all linear layers and are double-quantized in 4-bit NormalFloat during training. We set the alpha, rank, and dropout rates of the adapters to 16, 64, and 0.1, respectively. A 32-bit paged AdamW optimizer (Dettmers et al., 2024) is used with a batch size of 64 and a constant learning rate of 1e-4 (Wang et al., 2022; Wei et al., 2022a). Models are trained for 10 epochs with a maximum token length of 2,048 using NVIDIA A6000 (48GB VRAM) or A100 (80GB VRAM) GPUs. We use vLLM with greedy decoding for generation (Kwon et al., 2023). This setup is applied to all experiments unless otherwise specified.
Instructabiltiy evaluation. To evaluate whether RT models can handle user instructions, we assess their responses’ acceptability and quality. Additionally, we conduct pairwise assessments to measure their relative helpfulness compared to their IT counterparts.
  • Independent Assessment: We assess the acceptability of open-ended responses from RT models via human evaluation using the AlpacaEval test instructions (Li et al., 2023), which combine five evaluation datasets covering diverse instructions (Wang et al., 2023b; Köpf et al., 2023; Bai et al., 2022; Chiang et al., 2023; Geng et al., 2023). Evaluators are presented with an instruction and the corresponding model response and are asked to rate the responses by choosing one of three options: Acceptable (Excellent), Acceptable (Sufficient), or Not Acceptable. To address the limitations of human evaluators (Gudibande et al., 2024), we also conduct automatic fine-grained response quality evaluation using the JustEval benchmark (Lin et al., 2024), which utilizes GPT-4 as a judge. The evaluation interface, guidelines, and prompts can be found in Appendix A.
  • Pairwise Assessment: We conduct both human and automatic evaluations using the AlpacaEval test instructions. For the human evaluation, evaluators select the more helpful response or declare a tie between the responses provided by the RT model and its IT counterpart for the same instruction. For the automatic evaluation, we employ the GPT-4 judge from AlpacaEval (Li et al., 2023) and report length-controlled win rates (Dubois et al., 2024) of RT models against IT models. This judge exhibits a high Spearman correlation of 0.98 with human judgments in the Chatbot Arena (Chiang et al., 2024).
Core capabilities evaluation. To verify that RT models retain the foundational knowledge required for performing instructed tasks (Wang et al., 2023a), we evaluate their core capabilities using multiple benchmarks. The following benchmarks are considered: MMLU (Hendrycks et al., 2021) and OpenbookQA (Mihaylov et al., 2018) for knowledge, HellaSwag (Zellers et al., 2019) for commonsense, ARC (Clark et al., 2018) for reasoning, GSM8K (Cobbe et al., 2021) for mathematical reasoning, and PIQA (Bisk et al., 2020) for physical reasoning. The evaluation setup is detailed in Appendix A.
4.2 Results
RT models can appropriately respond to instructions. Figures 23 and 4 depict the human acceptability, response quality, and pairwise assessment results, respectively. The independent assessment results show that RT models—trained solely on responses—can generate appropriate answers to diverse instructions. A majority of the responses generated by RT models are rated as Acceptable, with many achieving the Excellent rating. They perform close to their IT counterparts in response quality evaluations, achieving similar scores across all metrics of the JustEval benchmark. In contrast, we observe that the base models rarely produce valid responses due to the absence of an established response distribution.
Additionally, the pairwise evaluation results indicate that while RT models can produce appropriate responses, their outputs are slightly less preferred compared to their IT counterparts. Our manual investigation of cases where IT model outputs are preferred over those of RT models finds that IT models often generate responses in a format better suited to the given instructions. These suggest that explicit instruction-response mapping helps models select a response format that is more appropriate from the learned distribution of responses. The examples of the model outputs are available in Appendix D. We further investigate how the training response distribution affects the model’s output in Appendices F and G.
In summary, our instructability evaluation results indicate that RT models—which are not explicitly trained to process instructions but rather learn solely the response distribution—can still provide helpful responses to diverse instructions. This finding suggests that instruction-processing ability is acquired during pre-training. [^4] We discuss the potential contamination concern regarding the IT datasets in Appendix H.
RT models mostly retain their core capabilities. Table 1 presents the core capabilities evaluation results. We find that RT models largely retain their core capabilities, with no notable degradation beyond the typical performance decline associated with fine-tuning (Wang et al., 2023a; Fu et al., 2024). Furthermore, across all tested tasks, RT and IT models exhibit similar performance, showing no substantial differences. These results suggest that RT models possess a knowledge base similar to that of IT models, which is required to perform the instructed tasks.
Instruction-processing ability is better internalized in larger LLMs. Table 2 shows the evaluation results for Gemma-2-2B and Gemma-2-9B RT models. While smaller models like Gemma-2-2B generate highly acceptable responses, larger models that have been pre-trained on substantially more tokens, such as Gemma-2-9B (2T vs. 8T tokens), exhibit higher human acceptance rates and smaller gaps with IT models. This suggests that larger models, with more extensive pre-training, can more effectively develop the ability to process instructions.
5 Rejecting Unsafe Instructions
In this section, we further investigate whether RT models can comprehend instructions and decide how to handle them. To test this, we introduce refusals for unsafe queries only through responses—thereby avoiding direct training for refusals via instruction-response mappings—and then evaluate whether the models can recognize and reject unsafe instructions.
@@ 5.1 Experimental Setup
Response tuning with refusals. We incorporate explanatory refusals for unsafe requests—responses that decline to fulfill requests and specify the underlying safety policy—into the response subsets of the Alpaca, Dolly, and LIMA datasets. These refusals are sourced from an existing LLM safety tuning dataset (unsafe query–refusal pairs) (Bianchi et al., 2024). To control their proportion, we add 100, 200, or 500 such examples into randomly sampled sets of 1,000 responses from the base RT datasets. [^5] We unify the size to match that of LIMA.
(f) Base LLM: Llama-3.1-8B (Dubey et al., 2024)
(g) Base LLM: Gemma-2-9B (Riviere et al., 2024)
Figure 7: GPT-4 response quality evaluation results for URIAL and URIAL-R. The test is conducted using the 1,000 test instructions from the JustEval benchmark, including the safety test set. The results show that URIAL-R achieves similar performance to URIAL across all metrics in both base models.
Evaluation. We measure the refusal rate (RR) of RT models for unsafe requests after training them solely on explanatory refusals. We then compare these rates to those of their IT counterparts, which are trained with the paired unsafe instructions. For evaluation, we use unsafe instructions from the following datasets: AdvBench (Zou et al., 2023), MaliciousInstruct (Huang et al., 2024), HarmBench (Mazeika et al., 2024), and XSTest (Röttger et al., 2024). To consider the exaggerated safety behaviors (excessive refusals), we also measure the False Refusal Rate (FRR) using benign queries from XSTest that are lexically similar to unsafe ones (Röttger et al., 2024). Existing methods for measuring RR often involve significant costs, require considerable human effort, or are unable to measure FRR. Therefore, we utilize an efficient and strong LLM—Llama-3.1-70B-Instruct (Dubey et al., 2024)—to classify whether a response is a refusal or compliance, i.e., whether the response declines to assist or attempts to provide the requested information. However, our initial inspection found that simply instructing the model to perform this classification yields poor performance due to these tricky cases:
@@
To universally and accurately detect refusals, we use a manually designed judge prompt that invokes chain-of-thought reasoning (Wei et al., 2022b) (see Appendix A). We validate this judge on 120 examples—60 compliance and 60 refusal responses, each with 30 of the edge cases described above—and it achieves 98.33% accuracy. We use this judge for all refusal evaluations.
@@ 5.2 Results
RT models can identify and reject unsafe requests. Figure 6 shows the evaluation results for the Gemma-2-9B model trained based on LIMA. The results show that RT models trained with refusals for unsafe queries exhibit substantially higher RR compared to those trained without refusals. It indicates that they are able to recognize and reject unsafe requests. We also find that their FRR falls within an acceptable range. Although they require more data, their refusal rates approach those of IT counterparts that are additionally supervised from mappings between unsafe queries and refusals. These results suggest that pre-training enables models to understand instructions and invoke their knowledge to process them appropriately.
@@ 6 In-Context Response Learning
We further validate our hypothesis in an in-context learning setting. To this end, we test whether untuned base LLMs can appropriately respond to user queries when provided only with demonstrations of responses.
Experimental setup. We remove instructions and the associated instruction-response mappings from URIAL (Lin et al., 2024), which consists of 4 instruction-response pairs including one pair of unsafe instruction and refusal. We refer to this new version as URIAL-R. We then evaluate it using two different base LLMs, Llama-3.1-8B, and Gemma-2-9B, with the JustEval benchmark. We employ greedy decoding with a maximum generation length of 2,048 tokens. We also evaluate the zero-shot template prompting baseline (Lin et al., 2024) to further investigate the effect of learning the response distribution. The prompts and details of the setup can be found in Appendix A.
Results. Figure 7 presents the evaluation results. The results show that using response demonstrations alone enables the base models to effectively handle both benign and unsafe instructions. Across all metrics, the scores of the model prompted with URIAL-R are similar to those of the model prompted with the original URIAL, which includes instructions. Additionally, while the zero-shot prompting baseline generates outputs relevant to the instructions, URIAL-R—which includes demonstrations of coherent responses—substantially outperforms it. These results further reinforce our earlier conclusion that the ability to handle instructions is inherent in pre-trained LLMs, and that establishing an adequate output distribution helps the models use them effectively.
@@ 7 Conclusion
We hypothesize that the pre-training stage enables LLMs to develop the ability to process instructions. To test this, we propose Response Tuning (RT), a method that removes instructions from IT and focuses solely on learning the response distribution. Our extensive experiments demonstrate that RT models—trained only on responses without paired instructions—effectively respond to a wide range of user queries. Moreover, we observe that they can reject unsafe requests by comprehending instructions and invoking learned conditions of refusal. These observations also hold in an in-context learning setting. Together, these results show that establishing an adequate output distribution alone can yield instructable models, supporting our hypothesis that pre-trained models possess the ability to process instructions.
Limitations
The primary aim of our work is to provide insights into how LLMs shift from pre-trained models to instructable assistants. In our study, RT serves as a tool—intentionally designed as a simplified ablation of IT—but as a standalone method, it exhibits limited performance compared to IT. Future work could build on the insights gained from RT to develop more effective and efficient tuning methods. Moreover, future research could explore the effect of pre-training on more nuanced aspects of instruction-following—such as mitigating sycophancy or social bias (Perez et al., 2023; Sharma et al., 2024).
@@
Our study involves human evaluations to evaluate instruction-following LLMs. The evaluators were hired in compliance with local laws and were paid appropriate compensation. The authors manually reviewed the LLM responses flagged by OpenAI moderation API and confirmed that these pose no harm to human evaluators. In addition, evaluators had the right to immediately stop the evaluation if they wished, and were encouraged to discuss any discomfort with the authors. While we publicly release the codes for safety evaluations, we decide not to release the refusal judge validation set to prevent potential misuse of unsafe or illegal information.
Reproducibility Statement
The training setup of the models including hyperparameters can be found in Section 4.1. We detail the evaluation setup in Appendix A.
@@ Acknowledgements
Seokhyun An is supported by the Korea Presidential Science Scholarship. This work was supported by grants from the Institute of Information & Communications Technology Planning & Evaluation (IITP), funded by the Ministry of Science and ICT (MSIT) of the Korean government: No. RS-2020-II201336 (Artificial Intelligence Graduate School Support Program at UNIST) and IITP-2024-RS-2024-00360227 (Leading Generative AI Human Resources Development). We also acknowledge the 2022 Research Fund (1.220140.01) of UNIST.
@@ A.2 Automatic Evaluations
  • MMLU (Hendrycks et al., 2021): We use the script from the open-instruct repository (Ivison et al., 2023) for the evaluation. [^8] https://github.com/allenai/open-instruct Exact-match accuracy is reported in a zero-shot setting.
@@
Refusal evaluation. We measure the refusal rates for unsafe instructions and false refusal rates for benign instructions using multiple evaluation datasets. For HarmBench (Mazeika et al., 2024), we report the average refusal rates for standard and contextual attack subsets. We use Llama-3.1-70B-Instruct (Dubey et al., 2024) with our judge prompt to classify refusals (see Table 7). This judge was validated using a set of 120 examples consisting of 60 compliance and 60 refusal responses, each containing 30 of the edge cases described in our experiment section. The edge cases are generated using GPT-4 and our internal jailbroken LLMs.
@@ Appendix B Experimental Setup
Response in-context learning. The simplified template of URIAL (Lin et al., 2024), URIAL-R, and zero-shot template prompt used for the evaluation can be found in Tables  45 and 6, respectively. We use urial1kv4 prompt in their official repository as a base URIAL prompt. [^9] https://github.com/Re-Align/URIAL/blob/main/urial_prompts/inst_1k_v4.txt.md Full version of URIAL-R prompt can be found in our code repository. The generation of the LLMs is truncated by the response marker of URIAL (‘‘‘).
@@ Appendix C Full Experimental Results
  • Refusal evaluation results: See Tables 11 and 12.
Appendix D Model Output Examples
@@
  • Llama-3.1-8B (Dubey et al., 2024): See Table 13.
  • Gemma-2-9B (Riviere et al., 2024): See Tables 14 and 15 (for the model trained with refusals).
@@
Appendix E Data Examples
Examples of explanatory refusals we used are presented in Table 18.
Appendix F RT without using IT Datasets
Theoretically, RT can also be performed using general texts that do not have paired prompts. In this experiment, we explore the RT using excerpted news articles.
Dataset. We randomly select 1,000 news passages from the CC-News dataset (Hamborg et al., 2017) and process them for use as RT training data. To establish a meaningful output distribution with news articles, we clean the excerpted sentences, enhance their readability, and adjust their tone using GPT-4o (OpenAI, 2023). The prompt and examples of the training data are provided in Tables 19 and 20.
Training and evaluation. Using this dataset, we fine-tune Llama-3.1-8B and assess its instructability with JustEval.
Results. Table 21 presents the evaluation results. The results indicate that the RT model, trained on excerpted news articles, can generate appropriate responses to various instructions, with a style reflecting the concise nature of its training data. Examples of the model’s outputs are shown in Table 22.
Appendix G Refining Response Distribution
Our experiments demonstrate that establishing a response distribution alone can make LLMs instructable, thanks to the inherent abilities acquired during pre-training. In this section, we explore whether refining the training response distribution can lead to improvements in user preference for the outputs. While previous works have shown that techniques such as feedback learning or fully regenerating IT dataset responses using aligned LLMs can enhance user preferences (Bai et al., 2022; Peng et al., 2023; Ivison et al., 2024), we investigate whether refining the response distributions in IT or RT data can yield similar benefits.
G.1 Experimental Setup
Response refinement. We focus on refining the fundamental characteristics of the response distribution, specifically examining three core elements that correlate with response quality: clarity, structure, and tone. To achieve this refinement, we use a strong instruction-following LLM, Llama-3.1-70B-Instruct (Dubey et al., 2024), with a manually crafted refinement prompt. The responses from the Alpaca, Dolly, and LIMA datasets are refined according to the following guidelines:
  • Clarity: Make the response easy to understand. It should be direct and to the point, avoiding complex language that might confuse the user.
The prompt and the example of refined data are available in Tables 23 and 24.
Evaluation. We conduct simulated preference evaluations using the GPT-4 judge of AlpacaEval (Li et al., 2023) and report length-controlled win rates (Dubois et al., 2024). To directly measure the gains from response distribution refinement, we compare both the IT and RT models trained using the refined datasets with those trained on the original datasets.
G.2 Results
Refined response distribution improves user preference. As shown in Table 25, the pairwise evaluation results indicate that both IT and RT models trained with the refined responses substantially outperform their counterparts trained on the original responses. The improvements are consistent across different base models and datasets, except for Alpaca. This exception may be due to Alpaca’s responses having limited room for improvement, as they are generated using GPT-4, a highly-aligned model using human feedback. These findings suggest that the training response distribution substantially contributes to the model’s ability to generate effective outputs, emphasizing the importance of the training response distribution in yielding instructable model. The examples of model output can be found in Table 26.
Appendix H Note on Potential Contamination Concerns
We use widely recognized open-weight models in our main experiments. However, since their pre-training datasets are not publicly disclosed, there may be concerns that these datasets contain text conforming to our chat template with special tokens (<|user|> and <|assistant|>). However, our main results demonstrate that base models cannot appropriately respond to instructions without an established response distribution, suggesting that such data were not meaningfully included in their pre-training datasets. Furthermore, we verify that our core observation holds for the open-data OLMo-7B model, whose authors have confirmed that such data are not included in its pre-training dataset (Hewitt et al., 2024). The results for this model are presented in Table 27.